114583:

A rectangular window in Jim's bedroom is 36 inches wide. The diagonal of the window is 44 inches. Which measurement is closest to the height of the window?

vietress:

You would use the pythagorean theorem to find the answer to this problem. \[a^2 + 36^2 = 44^2\]The first step is to solve for the exponents. 36 squared is 1296 44 squared is 1936 The new equation is: \[a^2 + 1296 = 1936\]Next, you subtract 1296 from both sides. This gives us: \[a^2 = 640\]The final step is to find the square root on both sides. \[\sqrt{a^2} = \sqrt{640}\]This gives us an answer of \[a \approx 25.29\] Hope this helped! If it did, please consider marking my answer as "Best Response"

114583:

Thanks.